OpenStudy (anonymous):

Write the equation for the line tangent to the circle x^2 + y^2 = 29 at the point (2, 5).

OpenStudy (anonymous):

\[ 2 x + 2 y y' =0\\ y '= -\frac x y\\ y'(2,5)= -\frac 2 5 \] The equation of the tangent is \[ y- 5 = -\frac 2 5 (x-2)\\ \]